OpenStudy (anonymous):

The bottom row of a stack of blocks contains11 blocks. The row above it contains 9 blocks. The next higher row contains 7 blocks. The rows continue in this pattern, and the top row contains a single block . How many blocks does the stack contain in all? a. 6 blocks c. 36 blocks b. 24 blocks d. 48 blocks

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

It is decreasing by 2 every row

OpenStudy (anonymous):

Each row subtracts 2 from the prior row, so that there are 11 - 2x = 1 decreases, or 5 reductions in total. There is one more row than reductions, so 6 rows. if you add the first and last row together, you get 12, same for the second and second to last, etc. there are therefor 6/2 = 3 such pairs, all summing to 12. Therefore, 3*12 = 36 total blocks.